OpenStudy (anonymous):

solve: 11z-33/9 divided by z-3/6

OpenStudy (anonymous):

\[((11z-33) \div 9) \div((z-3)\div6) = ((11z-33) \div 9) \times(6\div(z-3))\] \[= (6\times11\times(z-3))\div(9\times(z-3)) = 22\div3\]